mhchen:

Coding challenge for you guys: Create a javascript function that takes in integer, n, and prints (or console.log) all possible arrangements of the alphabet letters of length n. Example: if n = 1, then it'll just print a,b,c,d...z if n = 2, then it'll print, aa,ab,ac,ad....ba,bb,bc..zz if n = 3, it'll print aaa,aab,aac... and so on

mhchen:

Hint: Let _ represent a position then if n = 4, then it's ( _ _ _ ) + ( _ ) = ( _ _ ) + ( _ ) + ( _ ) = ( _ ) + ( _ ) + ( _ ) + ( _ ) and we know ( _ ) = a,b,c,d...